OpenStudy (anonymous):

-6p+7=3(2p-3)-4(-10+4p)

OpenStudy (anonymous):

solve for p

OpenStudy (anonymous):

-6p+7=6p-9+40-16p 4p=24 p=6

OpenStudy (anonymous):

p=6

OpenStudy (anonymous):

Use distributive law a(b + c) = ab + ac and a(b - c) = ab - ac -6p+7=3(2p-3)-4(-10+4p) -6p+7 = 6p - 9 + 40 - 16p -6p+7 = -10p + 31 -6p + 10p = 31 - 7 4p = 24 p = 24/4 p = 6

OpenStudy (anonymous):

thank you very much